A Bibliography of Publications of John Warner Backus

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

16 June 2026

Abstract

This bibliography records publications of John W. Backus.

Title word cross-reference

1539-1 [Ame97, ISO97a, ISO97b, ISO04a, ISO04b, ISO10]. 1539-2 [ISO94, ISO00]. 1539-3 [ISO99]. 1950s [Bac80a]. 1959 [Ano60]. 1960 [AGG60]. 1962 [Bus62]. 1966 [Ame66, AB66, Ame78b, Ame78c]. 1978 [Ame87a, AC87]. 1986 [SS86]. 1987 [Tur90]. 198X [AC87, Ame87b]. 1992 [AC92]. 1994 [ISO94]. 1997 [Ame97, ISO97a, ISO97b]. 1998 [ISO98a, ISO98b]. 1999 [ISO99]. 199x [Ame90].

2 [Spe83]. 2000 [ISO00]. 2007 [Bjø08]. 20th [Cip00]. 2DT [BARSW95]. 2DT-FP [BARSW95].

314 [Čul72b]. 367 [Čul72a].

4-1-4-1 [Int59].

50 [Gep66].

60 [BBG^{+}60a, BBG^{+}60b, BBG^{+}61, BBG^{+}1 , BBG^{+}62, BBG^{+}63c, BBG^{+}63a, BBG^{+}76, Čul72b, Čul72a, Dij62, Gep66, NBB^{+}60, NBB^{+}62, NBB^{+}63, Nau64, NBB^{+}65, NBB^{+}97]. 67 [Ame87a]. 67-R [Ame87a]. 6th [SS86].

701 [IBM53]. 704 [BA54]. 78-06-01 [Ame78d].

8X [Ame89].

AA [Bac75d]. Abstraction [HW87, HW95]. acceptor [KJ74]. ACM [Bac59, Ash87, Bac60, Wex81]. ACM-GAMM [Bac59, Bac60]. Address [Bac81b]. Afterword [Bac84a]. Ahead [DP82]. Aiken [BHW^{+}90]. Akademie [Gep66]. Akademie-Verlag [Gep66]. Algebra [Bac78a, Bac81c, Bac82b, Bac87a, Wil82b, Bac87b, Wil80]. Algebraic [Bac59, Bro85, DM83, Bac60]. ALGOL [BBG^{+}63c, Čul72b, Čul72a, NBB^{+}62, Gep66, BBG^{+}61, BBG^{+}1 , BBG^{+}62, BBG^{+}63c, BBG^{+}63a, BBG^{+}63b, BBG^{+}76, Dij62, NBB^{+}60, NBB^{+}62, NBB^{+}63, Nau64, NBB^{+}65, BBG^{+}60a, BBG^{+}60b, NBB^{+}97]. ALGOL-60 [BBG^{+}76, NBB^{+}65]. Algorithmic [BBG^{+}63c, BBG^{+}76, NBB^{+}60, NBB^{+}63, NBB^{+}97, dBvV81, BBG^{+}60a, BBG^{+}60b, BBG^{+}61, BBG^{+}1 , BBG^{+}62, BBG^{+}63a, BBG^{+}63b, Čul72b, Čul72a, Dij62, Gep66, Nau64, NBB^{+}65]. Algorithmische [BBG^{+}63c, Gep66, NBB^{+}62]. Algorithms [Cip00, Len93]. also [Čul72b, Čul72a]. Ambiguity [Can62]. America [Bac80a]. American [Ame78b, Ame78c, Ame90]. Amsterdam [dBvV81]. Analysis [TSF^{+}87]. Angeles [Wex81]. Annual [Len93]. ANS [Bac55]. ANSI [Ame78b, Ame78c, AC87, Ame66, Ame78b, Ame78c, Ame87a, Ame87b, AC92, Ame97]. ANSI/ISO/IEC [Ame97]. Anticipation [BHW^{+}90]. any [KJ74]. Appendix [BBG^{+}61, BBG^{+}62]. Application [Hud89, KJ74, PP24]. Applicative [Bac73d, McJ75, Bac73a]. approved [Ame78b, Ame78c]. April [Ame78b, Ame78c]. architectures [Bac82a]. Arithmetic [She59]. array [Hoe22]. arrogance [Bac79c, Che16]. ASA [Gor61]. Association [Čul72b, ABC^{+}54]. Atanasoff [BHW^{+}90]. August [Tur90]. Austin [Tur90]. Automatic [BBB^{+}56, Bac58, B^{+}96, Bau56, Int59, She59, BH54, BBB^{+}57]. Award [Ano94, Ash87, Dij78, Lew94, Ano10]. away [Bac79c, Che16].

B [Čul72b, Čul72a, Gep66]. Babbage [BHW^{+}90, BHW^{+}90, Ran91]. Backus [Bjø08, Čul72b, Čul72a, Gep66, Ing67, Kon73, Sap67, Chi80, Aik07, Ano94, Ano10, Ano22, Anoxx, Boo06, Bro85, CK07, Can62, CCH73, Che16, Dij78, DM83, EA06, IBMxx, Ing62, Joh66, KG62, Knu64, Kon73, KJ74, Loh07, MŞ11, McJ17, Nec22b, Nec22a, Nor86, Pag86, Pet65, PP24, Rad87, RD89, Ran91, Rob80, Roh68, Sla87a, Ste79, Y.77]. BACKUS-FP [Chi80]. Backus-type [Pag86]. Bad [Len93]. baker [Gor61]. Base [Ame97, ISO97a, ISO04a, ISO04b, ISO10, ISO97b]. based [Bac81b, Kon73].

Basic [AB66]. Bauer [Čul72b, Čul72a, Gep66]. Be [Bac78a, Bac87b, Bac87a]. Bericht [BBG^{+}63c, Gep66, NBB^{+}62]. Berkeley [Bad93]. Berlin [Gep66]. Best [Cip00]. between [Che16]. Biographies [Cor87, Wei88, BHW^{+}90]. BNF [BHW^{+}90]. Book [Čul72b, Čul72a]. bottleneck [Y.77]. Bowditch [BHW^{+}90]. breath [Bac79c, Che16]. Buchbesprechung [Gep66]. Bulletin [BBG^{+}63c].

C [Čul72b, Čul72a, Gep66]. calculus [Rev92]. Calif. [Wex81]. California [Sap67]. Can [Bac78a, Bac87b, Bac87a]. Centre [dBvV81]. Century [Cip00, MHR80]. character [ISO94, ISO00]. Charles [BHW^{+}90]. Church [Bac75d, McJ75, Rev92]. Clarification [ANS69, ANS71]. Class [Bac73b, Bac73c, Bac75c, Bac75d, Bac75a, Bac75b]. Closed [Bac73d, McJ75, Bac73a]. Code [SHS^{+}93]. Coding [BBB^{+}56, B^{+}96, Int59, She59, BBB^{+}57]. Collection [BBD^{+}81, MHR80]. combinatorial [Pos43]. combinators [Hoe22]. Combinatory [Hoe22]. Coming [Bac83]. Command [MA89]. Comments [BHW^{+}90]. Communications [Čul72b]. compilation [ISO99]. Compiler [She59]. Completeness [HW87, HWW90, HW95]. computable [Nec22b, Nec22a]. Computation [DP82]. Computer [Bac55, Čul72a, Knu03, Lap96, SL98, Bac81b, NKW70, SS86]. computer-oriented [NKW70]. Computing [ABC^{+}54, Bac83, Čul72b, MHR80, Bac82a]. concept [Bac79b, Bac81b]. Conception [Hud89]. Conditional [ISO99]. Conference [Ano60, Bac60, SS86, Wex81, Bac59]. Congress [Bac03]. Construction [Joh66, Kon73]. Contributions [Sam81]. Control [Bac55]. convenience [HW88]. crackles [Ste79]. Customs [PP24].

d’arbres [Rob80]. Data [BARSW95, Bus62, Cor87, ISO98b]. Days [Bac84b]. Debate [BHW^{+}90]. Dec [Bjø08]. December [SS86]. decision [Pos43]. Defining [Bac75d]. Definition [Joh66]. Definitions [Bac81c, Bac82b, Bac75b]. Delhi [SS86]. Denotational [HWWW85, MA89]. Design [AWW89, AWW94, BA54, Bac55]. Developer [Lew94, Loh07]. Development [Wil82b, Bac82a, Wil80]. Devices [Joh66, MŞ11]. Dictionary [Cor87]. Dies [Loh07]. different [Bac82a]. Dijkstra [Che16]. Dimensional [BARSW95]. Discoveries [SL98]. do [HW88]. document [BK93]. Does [BHW^{+}90]. dozen [Gor61]. Draft [Ame87c, Ame87b, Ame90, Ame87a, Ame89, ISO04a]. draw [HW88]. dynamically [AM91].

Early [Bac84b]. Edited [BBG^{+}63c]. Editor [Knu64, Čul72b, KG62]. Editors [Cip00]. EDPM [Int59, BBB^{+}56]. Edsger [Che16]. Efficient [DP82]. elimination [RD89]. Engineering [Lew94]. Enhanced [ISO98b]. Equations [Bac81c, Bac82b, LZ54, Bac79b]. Erinnerung [BBG^{+}63c]. Errors [AWW95]. ESA’93 [Len93]. Essays [MHR80]. Ethics [SHS^{+}93].

European [Len93]. evaluation [Rad87]. Evolution [Hud89]. Examples [BBD^{+}81, Gor61]. exception [ISO98a]. exhibit [Int57]. Expressible [Ing62]. expressions [BK93]. Expressive [HW87, HW95]. Extended [Ame90, Bac81c, Bac82b, EA06, AC92, Bac75d, Bac81b, BK93]. extending [Bac79b]. extension [Rev92].

F [Čul72b, Čul72a, Gep66]. facilities [ISO98b]. fame [Y.77]. feasibility [Pag86]. February [AGG60]. Fellow [Ano10]. Financial [PP24]. First [ABC^{+}54, Len93, Ash87, Pag86]. five [Int57]. FL [AWW89, AWW94, BWW86, BWW^{+}89, BWW90]. Floating [ISO98a]. Floating-point [ISO98a]. Form [Ing67, CCH73, Kon73, EA06, Ing62, Joh66, Knu64, Kon73, Pet65, Roh68]. Formal [L’E87, Luc81, Pos43]. Formatting [L’E87]. forms [Nec22b, Nec22a, PP24]. FORmula [IBM54, Anoxx]. FORTRAN [She59, Ame66, Ame87b, Int59, ISO04a, ISO04b, ISO10, AB66, Ame78b, Ame78a, Ame78c, Ame87a, AC92, BBB^{+}57, Bac58, Bac78b, Bac79a, Bac81a, Bac84b, B^{+}96, Int57, ISO94, Sla87a, Ame87c, ANS69, ANS71, Ame78d, AC87, Ame89, Ame90, Ame97, Ano93, BBB^{+}56, BH64, Bac80b, Bac98a, II91, ISO97a, ISO97b, ISO98a, ISO98b, ISO99, ISO00, Lew94, Loh07, Y.77]. Forum [SHS^{+}93]. Foundations [SS86]. Fourier [BHW^{+}90]. FP [Bac75a, Bad93, BARSW95, Chi80, HWWW85, HWW86, HW87, HWW90, HW95, Nor86, Pag86, Rad87, RD89, Wil81, Wil82a]. FP-systems [Rad87, RD89]. Francisco [Sap67]. free [Bac72a, Bac72b, Bac75a]. French [Rob80]. Full [HW87, HW95]. Function [Bac74, Bac81c, Bac81d, Bac82a, Bac85a, Bac85b]. Function-level [Bac82a]. Functional [AWW89, AWW94, Bac78a, Bac81c, Bac82b, Bac87a, BARSW95, Hud89, Rob87b, Wil82b, Wil82a, Bac75a, Bac79b, Bac87b, Bro85, Chi80, DM83, MŞ11, Pag86, Rob87a, Tur90, Wil80, Wil81]. fundamental [Bac81b]. funzionale [Chi80].

GAMM [Bac59, Bac60]. general [Pos43]. generalization [Nor86]. generating [Nec22b, Nec22a]. generator [KJ74]. German [BBG^{+}63c, Kon73, NBB^{+}62]. Germany [Len93]. Glossary [ABC^{+}54]. Good [HWW86]. Grammars [EA06, BK93, Kon73, Nec22b, Nec22a]. Grammatiken [Kon73]. Graphical [EA06, CCH73]. Great [Lap96, SL98]. greatly [Bac82a]. Green [Čul72b, Čul72a, Gep66]. guy [Bac79c, Che16].

H [Čul72b, Čul72a, Gep66]. handling [ISO98a]. he’d [Y.77]. held [Tur90]. Herausgegeben [BBG^{+}63c]. High [Sam81, Ano93]. Historical [Cor87]. History [All81, Bac78b, Bac79a, Boo06, BHW^{+}90, MHR80, Sam81, Wex81, Bac81a, Bac98a]. Honnef [Len93]. House [Ano60].

ibid [Čul72a]. IBM [IBM53, All81, BH54, Bac54, BA54, BBB^{+}56, Int59, Sam81, She59]. ideas [Ste79]. IEC [ISO04a, AC92, Ame97, ISO94, ISO97a, ISO97b, ISO98a, ISO98b, ISO99, ISO00, ISO04b, ISO10]. II [Bac58, Bac75b, Bac78b, Bac79a, Bac81a, Bac98a]. III [Bac75c, Bac78b, Bac79a, Bac81a, Bac98a]. Illinois [Rob87b, Rob87a]. Implementation [MA89]. Impressions [Bac80a]. Inc. [Ame78b, Ame78c]. India [SS86]. inference [AM91]. infinite [DM83]. Information [Ame89, Ame97, Ano60, ISO97a, ISO00, ISO04a, ISO04b, ISO10, AC87, II91, ISO94, ISO97b, ISO98a, ISO98b, ISO99]. Initial [ANS69]. Institute [Ame78b, Ame78c]. International [Ano60, Bac60, II91, ISO04a, dBvV81, Bac59]. Interpreter [MA89, Rob87a]. Interview [Sap67]. Introduction [BWW90, Mey90]. Invented [Sla87a]. inventor [Bac98b]. Investigation [PP24]. Invited [Bac81b]. ISO [Ame97, ISO04a, AC92]. ISO/IEC [ISO04a, AC92, ISO94, ISO97a, ISO97b, ISO98a, ISO98b, ISO99, ISO00, ISO04b, ISO10]. Italian [Chi80]. Items [Bac03]. IV [Bac75d].

J [Čul72b, Čul72a, Gep66]. J. [Čul72b, Čul72a]. John [Aik07, Ano94, Ano10, Ano22, Anoxx, Bjø08, Boo06, CK07, Che16, Dij78, IBMxx, Loh07, McJ17, Sap67, Sla87a, Ste79, Y.77]. Joseph [BHW^{+}90]. Journal [Čul72a]. June [Ano60, Wex81].

Katz [Čul72b, Čul72a, Gep66]. Kerner [Gep66]. Konstruktion [Kon73].

L [Čul72b, Čul72a, Gep66]. LALR [DP82]. lambda [Rev92]. lambda-calculus [Rev92]. langages [Rob80]. Language [Ame90, AWW89, AWW94, All81, BDE^{+}58, Bac60, BBG^{+}63c, Bac73d, BBG^{+}76, BWW^{+}89, BWW90, BBD^{+}81, BARSW95, NBB^{+}60, NBB^{+}63, NBB^{+}97, Ame87c, Ame78b, Ame78a, Ame78c, Ame87a, AC87, Ame89, AC92, Ame97, AM91, BBG^{+}60a, BBG^{+}60b, BBG^{+}61, BBG^{+}1 , BBG^{+}62, BBG^{+}63a, BBG^{+}63b, Bac73a, BWW86, Bro85, Čul72b, Čul72a, Dij62, DM83, Gep66, ISO97a, ISO97b, KG62, KJ74, MŞ11, NBB^{+}62, Nau64, NBB^{+}65, Pag86, Bac59, ISO04a, ISO04b, ISO10]. Languages [Bac73d, Bus62, Hud89, Ing62, Joh66, Knu03, Luc81, McJ75, Mey90, Ros66, Sam81, Ame97, Bac72a, Bac72b, Bac73b, Bac73c, Bac73a, Bac75c, Bac75d, Gor61, Hoe22, II91, ISO94, ISO97a, ISO97b, ISO98a, ISO98b, ISO99, ISO00, ISO04a, ISO04b, ISO10, Rob80, Wex81, dBvV81]. Lazy [Rad87]. Lecture [Dij78]. Lectures [Ash87]. length [ISO94, ISO00]. Letter [Knu64]. Letters [AGG60, Che16, KG62]. Level [Bac81c, Bac82b, Bac85a, Sam81, Bac74, Bac81d, Bac82a, Bac85b]. Lexing [Spe83]. Liberated [Bac78a, Bac87a, Bac87b]. Library [Bac03]. Life [BHW^{+}90]. line [HW88]. Linear [Bac81c, Bac82b, BHW^{+}90, Bac79b]. linked [Bac82a]. list [Rev92]. list-oriented [Rev92]. Lives [SL98]. logic [Hoe22]. Logical [TSF^{+}87]. Look [DP82]. Look-Ahead [DP82].

M [Čul72b, Čul72a, Gep66]. Machinery [Čul72b, ABC^{+}54]. make [Rad87]. Man [Sla87a]. manipulation [Rob80]. Manual [BBB^{+}56, BWW^{+}89, Bad93, Int59, BWW86]. March [Bjø08, Bus62]. Mathematical [IBM54, LZ54, dBvV81, Bac74, Bac81d]. Mathematik [Čul72b, Čul72a]. matrices [NKW70]. may [Bac82a]. McCarthy [Čul72b, Čul72a, Gep66]. MDN [Gep66]. mechanical [Gor61]. meeting [Tur90]. Memorandum [Bac72a]. Memoriam [Aik07]. memory [BBG^{+}63c]. method [Bac82a, KJ74]. Microprogramming [BHW^{+}90]. Minds [SL98]. Mitarb [Gep66]. ML [MA89]. Modified [BBG^{+}76]. Modula [Spe83]. Modula-2 [Spe83].

nach [Kon73]. Name [Cip00]. nano [MŞ11]. nano-devices [MŞ11]. National [Ame90]. Naur [Čul72b, Čul72a, EA06, Gep66, Knu64, PP24, Roh68, AGG60, BBG^{+}63c, CCH73, Kon73, Nec22b, Nec22a]. Naur-Form [Kon73]. Netherlands [dBvV81]. Neumann [Bac78a, Bac87b, Bac87a, Y.77]. No [BBG^{+}63c, Čul72a]. nondeterminism [Rad87]. Normal [Ing62, Joh66, Knu64, Pet65]. notation [KJ74]. note [Roh68]. Notes [Wil81, Wil82a, Bac73b, Bac73c, Bac75c, Bac75d, Bac75a, Bac75b, BHW^{+}90]. Numerische [Čul72b, Čul72a].

O [Gep66]. objects [Bac74, Bac81d, DM83]. occasion [dBvV81]. October [dBvV81]. Oh [Wei88]. Operator [Int59]. Optimization [Bac85a, Bac85b]. Oral [Boo06]. oriented [NKW70, Rev92]. other [BH54].

P [BBG^{+}63c, Čul72b, Čul72a, Gep66]. P. [BBG^{+}63c]. Pānini [Ing67]. Papers [Knu03, Lap96, Tur90]. Parallel [BARSW95]. Paris [Ano60]. parsing [Spe83]. Part [Ame97, ISO97a, ISO00, ISO04a, ISO04b, ISO10, ISO94, ISO97b, ISO99]. Parts [BWW^{+}89]. Pascal [L’E87]. Pathfinder [Ste79]. Performance [Ano93, Bac58]. Perlis [Čul72b, Čul72a, Gep66]. Personal [Bac80a]. Peter [AGG60]. pioneer [Ste79, BHW^{+}90]. Pioneers [Wei88]. point [ISO98a]. polynomially [Nec22b, Nec22a]. Portraits [Sla87b]. pp [Čul72b, Čul72a]. practical [Rad87]. Preis [Gep66]. Preliminary [AGG60, Int59]. Presence [AWW95]. presented [Gor61]. primer [Dij62]. Problem [Can62, TSF^{+}87, Pos43]. Proceedings [Bus62, Len93, SS86, Wex81, dBvV81]. Processing [Ano60, Bus62, Cor87]. Processor [All81]. processors [Gor61]. prof [dBvV81]. Program [Bac85a, LZ54, Bac79b, Bac81b, Bac85b, KJ74]. program-generator [KJ74]. programmazione [Chi80]. Programmer [BBB^{+}56]. Programming [Ame90, Ame97, ABC^{+}54, BDE^{+}58, Bac73a, Bac73d, Bac78a, Bac80a, Bac87a, BWW90, Bau56, BBD^{+}81, BARSW95, Hud89, ISO97a, ISO00, ISO04a, ISO04b, ISO10, Luc81, Mey90, Ran91, Rob87b, Ros66, Sam81, TSF^{+}87, Wil82a, Ame87c, Ame78b, Ame78a, Ame78c, Ame87a, AC87, Ame89, AC92, BH54, Bac58, Bac72b, Bac75a, Bac82a, Bac87b, Bro85, Chi80, DM83, II91, Pag86, Rob87a, Ste79, Tur90, Wex81, Wil81, BHW^{+}90, ISO94, ISO97b, ISO98a, ISO98b, ISO99, Bac72a]. Programs [AWW95, Bac78a, Bac81c, Bac82b, Bac87a, L’E87, Wil82b, Bac74, Bac81d, Bac87b, Wil80]. Progress [ANS69]. Project [Ame87a, AWW89, AWW94]. properties [Bac58, Bac75d]. Property [McJ75, Bac75d]. Proposal [BDE^{+}58]. Proposed [Ame87b, Ame90, Bac59, Ame87c, Ame87a, Bac60].

Queries [BHW^{+}90]. questions [Nec22b, Nec22a].

R [Ame87a]. radically [Bac82a]. Reasoning [Bac81c, Bac82b]. Receives [Lew94]. Recipient [Ano10]. recipients [Bac72a]. Recognition [Joh66]. Reconsidered [SHS^{+}93]. recursion [RD89]. Red [Bac73b, Bac73c, Bac75c, Bac75d]. Reduction [Bac72b, Bac72a]. reductions [Pos43]. Reference [BBB^{+}56]. Regular [Joh66, BK93]. Remembering [McJ17]. Report [ANS71, ABC^{+}54, AGG60, BBG^{+}60a, BBG^{+}60b, BBG^{+}61, BBG^{+}1 , BBG^{+}62, BBG^{+}76, NBB^{+}60, NBB^{+}63, NBB^{+}97, BBG^{+}63a, BBG^{+}63b, Čul72a, Dij62, Nau64, NBB^{+}65, BBG^{+}63c, Čul72b, Gep66, NBB^{+}62]. representation [CCH73, NKW70]. representations [Nec22b, Nec22a]. Research [Tur90]. Response [BHW^{+}90]. Restless [Bac98b]. retirement [dBvV81]. Rev [Bad93]. Review [Čul72b, Čul72a, Dij78]. Revised [BBG^{+}63a, BBG^{+}63b, NBB^{+}63, Nau64, NBB^{+}65, NBB^{+}97, Ame87c, Čul72a]. Revision [Ame78b, Ame78c, AC87, Ame87a]. Revolution [Bac83]. Rewrite [HWW90, HWWW85, HWW86]. Rome [Bus62]. Rosser [Bac75d, McJ75, Rev92]. Rules [HWW90, L’E87, HWWW85]. Russian [NBB^{+}65]. Rutishauser [Čul72b, Čul72a, Gep66].

S [Gep66]. S8 [AC87]. Sacrificing [HW88]. Safe [AWW95]. Samelson [Čul72b, Čul72a, Gep66]. San [Sap67]. satisfying [Rev92]. Schemes [RD89]. Science [BHW^{+}90, Lap96, Bac81b, SS86]. Scientists [SL98]. Second [ANS71]. Selected [Knu03]. Semantic [AWW95]. Semantics [Bac59, Bac60, Bac73d, Bac85a, Luc81, MA89, Bac73a, Bac85b, DM83, HWWW85]. September [Len93]. set [Gor61]. Sets [DP82]. SGML [BK93]. SIGPLAN [Wex81]. Silicon [Sla87b]. simplicity [HW88]. simplify [Bac82a]. sistema [Chi80]. Software [Bac82a, SS86]. solving [Bac79b]. Some [Bac80a, Nec22b, Nec22a]. sont [Rob80]. Specification [Gor61, Bro85]. Specifications [IBM54]. Spectrum [BBD^{+}81]. Speedcoding [BH54, Bac54, IBM53]. Sprache [BBG^{+}63c, Gep66, NBB^{+}62]. Stand [BHW^{+}90]. Standard [AB66, ISO04a, II91, Ame87c, Ame78b, Ame78a, Ame78c, Ame87a, AC87, Ame89, Ame90, AC92, MA89]. Standards [ANS69, ANS71]. Static [AM91]. still [Ste79]. Story [BHW^{+}90]. Strategies [HWW90, HWW86]. strings [ISO94, ISO00]. Structured [Bro85]. Study [BHW^{+}90]. Style [Bac78a, Bac87a, Wil82a, Bac87b, Wil81]. subcommittee [Gor61]. suggested [Ing67]. Supplement [BBG^{+}63c]. Symbolic [Bus62]. Symposium [Bus62, Len93, dBvV81]. Syntax [Bac59, Bac60, Ing62]. System [Bac54, BA54, Bac55, BBB^{+}56, B^{+}96, IBM53, IBM54, Int59, BBB^{+}57, Bac75a, Chi80, She59]. systèmes [Rob80]. Systems [Ame89, Can62, Ros66, AC87, BH54, Bac58, Rad87, RD89, Rob80].

takes [Bac79c, Che16]. teaching [Pag86]. Technical [Sam81]. Technique [AWW95, Ing62]. Techniques [Bac55]. Technology [All81, Ame97, II91, ISO94, ISO97a, ISO97b, ISO98a, ISO98b, ISO99, ISO00, ISO04a, ISO04b, ISO10, SS86]. Telegram [TSF^{+}87]. Terminology [ABC^{+}54]. Texas [Tur90]. Their [SL98, Gor61, Joh66, MA89]. theorem [Rev92]. theoretical [SS86]. Theory [Mey90]. Three [Bac75b]. Top [Cip00, Lew94]. topics [Tur90]. TR [ISO98a, ISO98b, Kon73, Kon73]. TR-grammars [Kon73]. TR-Grammatiken [Kon73]. Transformation [Bac85a, Bac85b]. Transforming [AWW95]. TRANSlating [IBM54]. Translation [Ing62, LZ54]. Translator [Anoxx, She59]. transpose [Bac75b]. tree [Rob80]. tree-manipulation [Rob80]. tribute [dBvV81]. Turanski [BBG^{+}63c, BBG^{+}63c]. Turing [Ash87, Dij78]. Tutorial [Rob87b]. Twentieth [MHR80]. Twenty [Int57, Ash87]. Twenty-five [Int57]. Two [BARSW95]. Two-Dimensional [BARSW95]. Type [BA54, IBM53, AM91, ISO98b, Pag86]. typed [AM91].

u [Gep66]. Übers [Gep66]. UML [EA06]. Unambiguity [BK93]. unclog [Y.77]. Unesco [Ano60]. University [Tur90]. unofficial [Ame78d]. Use [Bau56]. User [Bad93].

variable [Bac72a, Bac72b, Bac75a]. variable-free [Bac72b, Bac75a]. Varying [ISO00, ISO94]. Vauquois [Čul72b, Čul72a, Gep66]. VDL [Luc81]. Verlag [Gep66]. Version [Chi80]. Versione [Chi80]. vi [Gep66]. View [EA06]. Violations [PP24]. vol [Čul72b, Čul72a]. vs [Knu64].

W [Aik07, Che16, Čul72b, Čul72a, Gep66, Loh07]. Warner [Bjø08]. Wegstein [Čul72b, Čul72a, Gep66]. Where [HW88]. Whiggism [BHW^{+}90]. Whirlwind [BHW^{+}90, BHW^{+}90, LZ54]. Who [Sla87a]. Whose [Ing62]. Wide [BBD^{+}81]. Wijngaarden [Čul72b, Čul72a, dBvV81]. Wijngarden [Gep66]. William [BBG^{+}63c, BBG^{+}63c]. Woodger [Čul72b, Čul72a, Gep66]. Work [BHW^{+}90]. written [KJ74]. wrong [Bac81b].

X3 [Ame87a]. X3.10 [AB66]. X3.10-1966 [AB66]. X3.198 [Ame90, AC92]. X3.198-1992 [AC92]. X3.198-199x [Ame90]. x3.4 [Gor61]. X3.9 [Ame78b, Ame78c, AC87, Ame66, Ame78b, Ame78c, Ame87a, AC87, Ame87b]. X3.9-1966 [Ame66, Ame78b, Ame78c]. X3.9-1978 [Ame78b, Ame78c, Ame87a, AC87]. X3.9-198X [AC87, Ame87b]. X3J3 [Ame78d]. X3J3/ [Ame78d].

years [Ash87, Int57, Ste79].

Zur [BBG^{+}63c]. Zurich [Bac59, Bac60].

References

ANSI:1966:ANS

[AB66] American National Standards Institute and Business Equipment Manufacturers Association (U.S.). American National Standard Basic FORTRAN: X3.10-1966. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1966. 32 pp.

Adams:1954:RAC

[ABC^{+}54] C. W. Adams, J. W. Backus, J. W. Carr, III, Grace Murray Hopper, Chair, R. F. Osborn, G. W. Patterson, J. Svigals, and J. Wegstein. Report to the Association for Computing Machinery: First glossary of programming terminology. Report, ACM Press, New York, NY 10036, USA, June 2, 1954. ii + 25 pp. URL https:/

/archive.computerhistory.org/resources/text/Knuth_Don_
X4100/PDF_index/k-8-pdf/k-8-u2741-2-ACM-Glossary.pdf.

ANSI:1987:ANSd

[AC87] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information systems: programming language Fortran: S8 (X3.9-198X): Revision of ANSI X3.9-1978. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1987. various pp.

ANSI:1992:ANSc

[AC92] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for programming language, FORTRAN — extended: ANSI X3.198-1992: ISO/IEC 1539: 1991 (E). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, September 21, 1992. xix + 369 pp.

Arden:1960:LPR

[AGG60] B. Arden, B. Galler, and R. Graham. Letters: Preliminary report (February 4, 1960) of Peter Naur. Communications of the ACM, 3 (6):A13, June 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [NBB^{+}60].

Aiken:2007:MJW

[Aik07] Alex Aiken. In memoriam: John W. Backus, 1924–2007. ;login: the USENIX Association newsletter, 32(3):68–69, June 2007. CODEN LOGNEM. ISSN 1044-6397. URL https://

www.usenix.org/publications/login/june-2007-volume-32-
number-3/memoriam-john-w-backus-1924%E2%80%932007.

Allen:1981:HLP

[All81] F. E. Allen. The history of language processor technology in IBM. IBM Journal of Research and Development, 25(5):535–548, September 1981. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Aiken:1991:STI

[AM91] Alex Aiken and Brian Murphy. Static type inference in a dynamically typed language. In David S. Wise, editor, Conference record of the eighteenth annual ACM Symposium on Principles of Programming Languages: Papers presented at the symposium, Orlando, Florida, January 21–23, 1991, POPL 91, pages 279–290. ACM Press, New York, NY 10036, USA, 1991. ISBN 0-89791-419-8 (paperback). LCCN QA76.7 .A15 1991.

ANSI:1966:AF

[Ame66] American National Standards Institute. ANSI Fortran X3.9-1966. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1966. ?? pp. Approved March 7, 1966 (also known as Fortran 66). See also subsequent clarifications [ANS69] and [ANS71].

ANSI:1978:ANSc

[Ame78a] American National Standards Institute. American National Standard programming language, FORTRAN. American National Standard; ANSI X3.9-1978 CSA standard; Z243.18-1980 American National Standards Institute. American National Standard; ANSI X3.9-1978. Canadian Standard Association. CSA standard; Z243.18-1980. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, revised edition, 1978. 438 pp. Available on the World-Wide Web at http://observer.

gsfc.nasa.gov/iteams/doc/ansi_f77.ps, http://observer.
gsfc.nasa.gov/iteams/doc/f77.doc, http://observer.gsfc.

nasa.gov/iteams/doc/f77_cov.pdf, and http://observer.
gsfc.nasa.gov/iteams/doc/f77_doc.pdf.

ANSI:1978:ANSb

[Ame78b] American National Standards Institute. American National Standard programming language FORTRAN: approved April 3, 1978, American National Standards Institute, Inc.: ANSI X3.9-1978. Revision of ANSI X3.9-1966. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, revised edition, 1978. 438 pp.

ANSI:1978:ANSd

[Ame78c] American National Standards Institute. American National Standard programming language FORTRAN: approved April 3, 1978, American National Standards Institute, Inc.: ANSI X3.9-1978. Revision of ANSI X3.9-1966. World-Wide Web docu-

ment., 1978. URL http://www.fortran.com/fortran/F77_std/

rjcnf0001.html. This is an online hypertext of the Fortran 77 Standard.

ANSI:1978:FUX

[Ame78d] American National Standards Institute.Committee on Computers X3 Information Processing. Fortran 77: unofficial, X3J3/ 90.5, (78-06-01). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1978. 371 pp.

ANSI:1987:ANSc

[Ame87a] American National Standards Institute. American National Standard programming language, FORTRAN: draft proposed: revision of ANSI X3.9-1978. X3 Project 67-R. Global Engineering Documents, Washington, DC, USA, revised edition, 1987. various pp.

ANSI:1987:DPA

[Ame87b] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed ANSI Fortran X3.9-198x, September 18, 1987.

ANSI-X3J3:1987:DPR

[Ame87c] American National Standards Institute.Technical Committee X3J3. Draft proposed revised American National Standard programming language Fortran. X3 Secretariat/CBEMA, Washington, DC, USA, version 104 edition, 1987. various pp.

ANSI:1989:ANSc

[Ame89] American National Standards Institute. American National Standard for Information Systems programming and language: Fortran 8X draft, volume 8(4) of Fortran Forum. ACM Press, New York, NY 10036, USA, May 1989. various pp. Revision of X3.9-1978.

ANSI:1990:DPA

[Ame90] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed American National Standard Programming Language Fortran Extended X3.198-199x, September 24, 1990.

ANSI:1997:AII

[Ame97] American National Standards Institute. ANSI/ISO/IEC 1539-1:1997: Information technology — Programming languages — Fortran — Part 1: Base language. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1997. ISBN ???? ???? pp. LCCN ???? US$18.00. URL http:

//webstore.ansi.org/ansidocstore/product.asp?sku=ANSI%
2FISO%2FIEC+1539%2D1%3A1997;   http://www.fortran.com/
fortran/iso1539.html.

Anonymous:1960:ICI

[Ano60] Anonymous, editor. International Conference on Information Processing, Unesco House, Paris, 15–20 June 1959. UNESCO and R. Oldenbourg and Butterworths, Paris, France and Munich, Germany and London, UK, 1960. URL https://archive.

computerhistory.org/resources/access/text/2019/02/102785380-
05-01-acc.pdf;   https://softwarepreservation.computerhistory.
org/ALGOL/paper/AP_Symposium-ICIP-1959.pdf.

Anonymous:1993:HPF

[Ano93] Anonymous. High Performance Fortran. Scientific Programming, 2(1-2):1–170, Spring–Summer 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:1994:AJB

[Ano94] Anonymous. Award for John Backus. ACM SIGPLAN FORTRAN Forum, 13(1):17, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic). URL https://dl.acm.org/doi/

pdf/10.1145/191559.1041030.

Anonymous:2010:JBF

[Ano10] Anonymous. John Backus: 1997 Fellow Award recipient.

Web document, 2010. URL https://web.archive.org/web/
20100709005030/http://www.computerhistory.org/fellowawards/

hall/bios/John,Backus/. For his development of FORTRAN, contributions to computer systems theory and software project management.

Anonymous:2022:JB

[Ano22] Anonymous. John Backus. Web site, 2022. URL http://

www.columbia.edu/cu/computinghistory/backus.html. Con-

tains links to a video of an address by John Backus to the Los Alamos International History of Computing Conference, June 1976.

Anonymous:20xx:JBF

[Anoxx] Anonymous. John Backus: Formula translator. Web doc-

ument., 20xx. URL https://lemelson.mit.edu/resources/
john-backus.

ANSI:1969:CFS

[ANS69] ANSI Subcommittee X3J3. Clarification of Fortran standards— initial progress. Communications of the ACM, 12:289–294, 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Ame66, ANS71].

ANSI:1971:CFS

[ANS71] ANSI Subcommittee X3J3. Clarification of Fortran standards— second report. Communications of the ACM, 14:628–642, 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Ame66, ANS69].

Ashenhurst:1987:ATA

[Ash87] Robert L. Ashenhurst, editor. ACM Turing Award Lectures: the first twenty years, 1966–1985. ACM Press anthology series. ACM Press and Addison-Wesley, New York, NY 10036, USA and Reading, MA, USA, 1987. ISBN 0-201-07794-9. xviii + 483 pp. LCCN QA76.24 .A33 1987.

Aiken:1989:FPD

[AWW89] Alexander Aiken, John H. Williams, and Edward L. Wimmers. The FL Project: The design of a functional language. Technical report, Computer Science Division, University of California, Berkeley and IBM Almaden Research Center, Berkeley, CA, USA and San Jose, CA, USA, 1989. ?? pp. URL https://theory.

stanford.edu/~aiken/publications/trs/FLProject.pdf. See

revision [AWW94].

Aiken:1994:FPD

[AWW94] Alexander Aiken, John H. Williams, and Edward L. Wimmers. The FL Project: The design of a functional language. Report, Computer Science Division, University of California, Berkeley and IBM Almaden Research Center, Berkeley, CA, USA and San Jose, CA, USA, December 11, 1994. 40 pp. URL https://theory.

stanford.edu/~aiken/publications/trs/FLProject.pdf.

Aiken:1995:SST

[AWW95] Alexander Aiken, John H. Williams, and Edward L. Wimmers. Safe: a semantic technique for transforming programs in the presence of errors. ACM Transactions on Programming Languages and Systems, 17(1):63–84, January 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0164-0925/201002.html.

Backus:1996:FAC

[B^{+}96] J. W. Backus et al. The FORTRAN automatic coding system. In Laplante [Lap96], chapter 2.1, pages 62–?? ISBN 0-314-06365-X (paperback), 0-7803-1112-4 (hardcover). LCCN QA76 .G686 1996.

URL http://bit.csc.lsu.edu/~chen/GreatPapers.html. Con-

tains a reprint of the first published paper on Fortran [BBB^{+}57].

Backus:1954:SDI

[BA54] John W. Backus and Gene Amdahl. The system design of the IBM Type 704. Talk given at an [unknown] ACM meeting,

1954. URL https://softwarepreservation.computerhistory.
org/FORTRAN/#By_others.

Backus:1954:ISS

[Bac54] J. W. Backus. The IBM 701 Speedcoding system. Journal of the ACM, 1(1):4–6, January 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://archive.

computerhistory.org/resources/text/Fortran/102663108.
05.01.acc.pdf.

Backus:1955:CSD

[Bac55] John W. Backus. Computer system design and ANS control techniques. Report, IBM Corporation, New York, NY, USA, October 26, 1955. 24 pp. URL https://mark.

people.clemson.edu/Backus-Computer_System_Design_and_
ANS_Control_Techniques-1955_10_26.pdf.

Backus:1958:APP

[Bac58] J. W. Backus. Automatic programming: properties and performance of FORTRAN systems I and II. In ????, editor, Proceedings of the Symposium on the Mechanisation of Thought Processes, Teddington, Middlesex, England, The National Physical Laboratory, November 1958, pages 232–255. Her Majesty’s Stationery Office, London, UK, 1958. URL https://archive.

computerhistory.org/resources/text/Fortran/102663114.
05.01.acc.pdf.

Backus:1959:SSP

[Bac59] J. W. Backus. The syntax and semantics of the proposed International Algebraic Language of the Zurich ACM-GAMM Conference. In ????, editor, Information Processing: Proceedings of the International Conference on Information Processing, pages 125– 132. R. Oldenbourg, München, Germany, 1959. URL https:/

/softwarepreservation.computerhistory.org/ALGOL/paper/
Backus-ICIP-1959.pdf;   https://unesdoc.unesco.org/ark:
/48223/pf0000153719.

Backus:1960:SSP

[Bac60] J. W. Backus. The syntax and semantics of the Proposed International Algebraic language of the Zurich ACM-GAMM conference. In Anonymous [Ano60], pages 125–

132. URL https://archive.computerhistory.org/resources/
access/text/2019/02/102785380-05-01-acc.pdf;   https:/
/softwarepreservation.computerhistory.org/ALGOL/paper/
AP_Symposium-ICIP-1959.pdf.

Backus:1972:MRR

[Bac72a] John Backus. Memorandum to recipients of Reduction languages and variable free programming. Report ???, IBM Research Laboratory, San Jose, CA, USA, April 19, 1972. 3 pp.

Backus:1972:RLV

[Bac72b] John Backus. Reduction languages and variable-free programming. Report RJ 1010, IBM Research Laboratory, San Jose, CA, USA, April 7, 1972. 111 pp. One-page errata issued on 8 May 1973.

Backus:1973:PLSa

[Bac73a] John Backus. Programming language semantics and closed applicative languages. Report RJ 1245, IBM Research Laboratory, San Jose, CA, USA, July 5, 1973. 38 pp. One-page errata issued on 17 July 1973 an 8 August 1973.

Backus:1973:CNRa

[Bac73b] John W. Backus. Class notes: Red languages. Notes., November 14, 1973.

Backus:1973:CNRb

[Bac73c] John W. Backus. Class notes: Red languages. Notes., November 21, 1973.

Backus:1973:PLSb

[Bac73d] John W. Backus. Programming language semantics and closed applicative languages. In Patrick C. Fischer and Jeffrey D. Ullman, editors, Conference Record of the ACM Symposium on Principles of Programming Languages, Boston, Massachusetts, USA, October 1973, pages 71–86. ACM Press, New York, NY 10036, USA, 1973.

Backus:1974:FLP

[Bac74] John Backus. Function level programs as mathematical objects. ACM SIGPLAN Notices, 9(4):1–10, April 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Backus:1975:CNVa

[Bac75a] John W. Backus. Class notes, I: A variable-free functional programming system, FP. Notes., February 19, 1975.

Backus:1975:CNVb

[Bac75b] John W. Backus. Class notes, II: Three definitions for transpose. Notes., March 7, 1975.

Backus:1975:CNIa

[Bac75c] John W. Backus. Class notes, III: Red languages. Notes., April 30, 1975.

Backus:1975:CNIb

[Bac75d] John W. Backus. Class notes, IV: Defining AA; properties of Red languages (extended Church–Rosser property). Notes., May 7, 1975.

Backus:1978:CPL

[Bac78a] John Backus. Can programming be liberated from the von Neumann style? A functional style and its algebra of programs. Communications of the ACM, 21(8):613–641, August 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317

(electronic). URL http://www.stanford.edu/class/cs242/

readings/backus.pdf. Reproduced in Selected Reprints on Dataflow and Reduction Architectures, ed. S. S. Thakkar, IEEE, 1987, pp. 215–243, and in ACM Turing Award Lectures: The First Twenty Years, ACM Press, 1987, pp. 63–130.

Backus:1978:HFI

[Bac78b] John Backus. The history of FORTRAN I, II, and III. ACM SIG-PLAN Notices, 13(8):165–180, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Backus:1979:HFI

[Bac79a] John Backus. The history of FORTRAN I, II, and III. Annals of the History of Computing, 1(1):21–37, July/September 1979. CODEN AHCOE5. ISSN 0164-1239. URL http://dlib.

computer.org/an/books/an1979/pdf/a1021.pdf;   http://
www.computer.org/annals/an1979/a1021abs.htm.

Backus:1979:ECP

[Bac79b] John W. Backus. On extending the concept of program and solving linear functional equations. Draft paper distributed at Summer Workshop on Programming Methodology, University of California at Santa Cruz, Santa Cruz, CA, USA. A revised version was later published [Bac81c]., August 1979.

Backus:1979:GAT

[Bac79c] John W. Backus. This guy’s arrogance takes your breath away. Letters between John W. Backus and Edsger W. Dijkstra., 1979.

URL https://medium.com/@acidflask/this-guys-arrogance-
takes-your-breath-away-5b903624ca5f.

Backus:1980:PAS

[Bac80a] John Backus. Programming in America in the 1950s: Some personal impressions. In Metropolis et al. [MHR80], pages 125– 135. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5

.I63 1976. URL https://www.sciencedirect.com/science/
article/pii/B9780124916500500174. Original versions of these

papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Backus:1980:F

[Bac80b] John W. Backus. Fortran, 1980.

Backus:1981:HFI

[Bac81a] J. Backus. The history of FORTRAN I, II, and III. In Wexelblat [Wex81], pages 25–74. ISBN 0-12-745040-8. LCCN QA76.7 .H56

1978. URL https://dl.acm.org/doi/book/10.1145/800025.

Backus:1981:CSB

[Bac81b] John Backus. Is computer science based on the wrong fundamental concept of ‘program’? An extended concept (invited address). In de Bakker and van Vliet [dBvV81], pages 133–165. ISBN 0-444-86285-4. LCCN QA76.7 .I575 1981. URL https://ir.cwi.nl/

pub/34328/34328D.pdf.

Backus:1981:AFP

[Bac81c] John W. Backus. The algebra of functional programs: Function level reasoning, linear equations, and extended definitions. In Josep Díaz and Isidro Ramos, editors, Formalization of Programming Concepts, International Colloquium, Peñíscola, Spain, April 19–25, 1981, Proceedings, volume 107 of Lecture Notes in Computer Science, pages 1–43. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 3-540-10699-5.

Backus:1981:FLP

[Bac81d] John W. Backus. Function level programs as mathematical objects. In Arvind and Jack B. Dennis, editors, Proceedings of the 1981 conference on Functional programming languages and computer architecture, FPCA 1981, Wentworth, New Hampshire, USA, October 1981, FPCA ’81, pages 1–10. ACM Press, New York, NY 10036, USA, 1981.

Backus:1982:SFL

[Bac82a] J. Backus. Software I: Function-level computing: a new programming method, linked to radically different architectures, may greatly simplify software development. IEEE Spectrum, 19(8):22– 27, August 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Backus:1982:AFP

[Bac82b] John Backus. The algebra of functional programs: Functional level reasoning, linear equations, and extended definitions. Report RJ 3555, IBM Research Laboratory, 5600 Cottle Road, San Jose, CA 95193, USA, July 23, 1982. 43 pp. This appears to be a reprint of [Bac81c].

Backus:1983:CRC

[Bac83] John W. Backus. The coming revolution in computing. Report RJ 3994, IBM Research Laboratory, San Jose, CA 95193, USA, August 23, 1983.

Backus:1984:A

[Bac84a] John Backus. Afterword. Annals of the History of Computing, 6 (1):26–27, January/March 1984. CODEN AHCOE5. ISSN 0164-

1239. URL http://dlib.computer.org/an/books/an1984/pdf/
a1015.pdf;   http://www.computer.org/annals/an1984/a1015abs.
htm.

Backus:1984:EDF

[Bac84b] John Backus. Early days of FORTRAN. Annals of the History of Computing, 6(1):15, January/March 1984. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1984/pdf/a1015.pdf;   http://www.computer.org/
annals/an1984/a1015abs.htm.

Backus:1985:FLSa

[Bac85a] John Backus. From function level semantics to program transformation and optimization. Report RJ 4567 (40935), IBM Research Laboratory, San Jose, CA 95193, USA, January 8, 1985. 35 pp.

Backus:1985:FLSb

[Bac85b] John Backus. From function level semantics to program transformation and optimization. In Mathematical foundations of software development, Vol. 1 (Berlin, 1985), volume 185 of Lecture Notes in Comput. Sci., pages 60–91. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 3-540-15198-2.

Backus:1987:CPLb

[Bac87a] John Backus. Can programming be liberated from the von Neumann style? A functional style and its algebra of programs. In Ashenhurst [Ash87], pages 63–130. ISBN 0-201-07794-9. LCCN QA76.24 .A33 1987. ACM Turing Award lecture.

Backus:1987:CPLa

[Bac87b] John W. Backus. Can programming be liberated from the von Neumann style? A functional style and its algebra of programs. Report RJ 2234, IBM Research Laboratory, San Jose, CA, USA, 1987. 98 pp. Draft of ACM Turing Award lecture.

Backus:1998:HFI

[Bac98a] J. Backus. The history of Fortran I, II, and III. IEEE Annals of the History of Computing, 20(4):68–78, October/December 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL https://ieeexplore.ieee.org/iel4/85/15706/
00728232.pdf.

Backus:1998:RI

[Bac98b] John Backus. Restless inventor. In Shasha and Lazere [SL98], page ?? ISBN 0-387-98269-8. LCCN QA76.2.A2 S53 1998. US$16.00.

Backus:2003:ILC

[Bac03] John Backus. Items for the Library of Congress. Web document,

2003. URL https://softwarepreservation.computerhistory.
org/FORTRAN/Backus%20-%20LOC%20-%20catalogue%20of%20papers.
pdf.

Baden:1993:BFU

[Bad93] Scott Baden. Berkeley FP User’s Manual, Rev. 4.1. Department of Computer Science, University of California, Berkeley, Berkeley, CA, USA, April 8, 1993. 39 pp. URL https://maibriz.de/unix/

ultrix/_root/fp.pdf.

Ben-Asher:1995:FPF

[BARSW95] Yosi Ben-Asher, Gudula Runger, Assaf Schuster, and Reinhard Wilhelm. 2DT-FP: a parallel functional programming language on two-dimensional data. International Journal of Parallel Programming, 23(5):389–422, October 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Bauer:1956:UAP

[Bau56] Walter F. Bauer. Use of automatic programming. Computers and Automation, 5(11):6–11, November 1956. CODEN CPAUAJ. ISSN 0010-4795, 0887-4549. The online edition of the Oxford English Dictionary cites this as the earliest mention of the name FORTRAN, with the quote: “John Backus’ group at IBM has prepared FORTRAN (FORmula TRANslation) for the IBM-704 computer. FORTRAN will translate into computer language a program written very close [sic] the language of the mathematician or scientist.” The OED citation references issue “Nov. 9/2”; that refers to page 9, column 2, of the article. There are no further references to Fortran in this journal up to at least 1962.

Backus:1956:FAC

[BBB^{+}56] J. W. Backus, R. J. Beeber, S. Best, R. Goldberg, H. L. Herrick, R. A. Hughes, L. B. Mitchell, R. A. Nelson, R. Nutt, D. Sayre, P. B. Sheridan, H. Stern, and I. Ziller. The Fortran Automatic Coding System for the IBM 704 EDPM: Programmer’s Reference Manual. Applied Science Division and Programming Research Department, International Business Machines Corporation, New York, NY, USA, October 15, 1956. 51 pp.

LCCN ???? URL https://archive.computerhistory.org/
resources/text/Fortran/102649787.05.01.acc.pdf.

Backus:1957:FAC

[BBB^{+}57] J. W. Backus, R. J. Beeber, S. Best, R. Goldberg, L. M. Haibt, H. L. Herrick, R. A. Nelson, D. Sayre, P. B. Sheridan, H. Stern, I. Ziller, R. A. Hughes, and R. Nutt. The FOR-TRAN automatic coding system. In Proceedings of the Western Joint Computer Conference, February 26–28, 1957, Los Angeles, CA, USA, pages 188–198. Institute of Radio Engineers, 1 East 79th Street, New York 21, NY, USA, 1957. ISSN 0449-1173. LCCN TK7885.A1 J6. URL http://bitsavers.org/

pdf/ibm/704/FORTRAN_paper_1957.pdf;   https://archive.
computerhistory.org/resources/text/Fortran/102663113.
05.01.acc.pdf;   https://softwarepreservation.computerhistory.
org/FORTRAN/#By_FORTRAN_project_members. The online edi-

tion of the Oxford English Dictionary cites this as the second earliest mention of the name FORTRAN, with the extract “The programmer attended a one-day course on FORTRAN and … then programmed the job in four hours using 47 FORTRAN statements.”.

Bauer:1981:PWS

[BBD^{+}81] F. L. Bauer, M. Broy, W. Dosch, R. Gnatz, B. Krieg-Brückner, A. Laut, M. Luckmann, T. Matzner, B. Möller, H. Partsch, P. Pepper, K. Samelson, R. Steinbrüggen, M. Wirsing, and H. Wössner. Programming in a wide spectrum language: a collection of examples. Science of Computer Programming, 1(1–2):73–114, October 1981. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Backus:1960:RALa

[BBG^{+}60a] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, Editor, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language Algol 60. Numerische Mathematik, 2(1):106–136, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

https://link.springer.com/article/10.1007/BF01386216.

Backus:1960:RALb

[BBG^{+}60b] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, Editor, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language Algol 60. Acta Polytechnica Scandinavica: Mathematics and Computing Machinery Series, (5):1–39, ???? 1960. CO-

DEN APSMAT.  ISSN 0355-2713.  URL https://web.

archive.org/web/20191029151524id_/http://web.eah-jena.
de/%7Ekleine/history/languages/Algol60-Naur.pdf.

Backus:1961:ARA

[BBG^{+}61] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, H. Rutishauser, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Appendix: Report on the algorithmic language ALGOL 60. Annual Review in Automatic Programming, 2:351–390, January 1961. CO-DEN ARVAAM. ISSN 0066-4138 (print), 1878-545x (electronic).

URL https://www.sciencedirect.com/science/article/pii/
S0066413861800165.

Backus:1962:RAL

[BBG^{+}62] John W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur (Editor), A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. Van Wijngaarden, and M. Woodger. Appendix: Report on the algorithmic language AL-GOL 60, pages 74–114. Volume 2 of APIC studies in data processing [Dij62], 1962. ISBN 0-12-216250-1. ISSN 0067-2483. LCCN

QA76.73.A24. URL https://dn710009.ca.archive.org/0/
items/a_primer_of_algol_60_programming/algol60_o.pdf.

Backus:1963:RRAa

[BBG^{+}63a] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language AL-GOL 60. The Computer Journal, 5(4):349–367, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/5/
4/349.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/050349.sgm.abs.html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/349.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/tiff/351.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/352.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/353.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/354.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/tiff/356.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/357.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/358.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/359.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/tiff/361.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/362.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/363.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/364.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_05/Issue_04/tiff/365.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/tiff/366.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/367.tif;   https://datamuseum.
dk/bits/30002299;   https://standardpascaline.org/Algol60-
RevisedReport.pdf;   https://web.archive.org/web/20200214113425id_
/http://web.eah-jena.de/~kleine/history/languages/Algol60-
RevisedReport.pdf.

Backus:1963:RRAb

[BBG^{+}63b] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL 60. Numerische Mathematik, 4:420–453, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/BF01386340. Edited by Peter

Naur. Dedicated to the memory of William Turanski.

Backus:1963:BAS

[BBG^{+}63c] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, Heinz Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Bericht über die Algorithmische Sprache ALGOL 60 (ALGOL Bulletin Supplement No. 3). Zur Erinnerung an William Turanski. Herausgegeben von P. Naur. (German) [Report on the algorithmic language ALGOL 60 (ALGOL Bulletin supplement no. 3). In memory of William Turanski. Edited by P. Naur]. Elektronische Datenverarbeitung, 2(??):1–24, 1963. ISSN 0374-3012.

Backus:1976:MRA

[BBG^{+}76] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, M. Woodger, R. M. De Morgan, I. D. Hill, and B. A. Wichmann. Modified report on the algorithmic language ALGOL-60. The Computer Journal, 19(4):364–379, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/4/364.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_04/tiff/364.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/
Issue_04/tiff/365.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_19/Issue_04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/367.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/368.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_19/Issue_04/tiff/369.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_04/tiff/370.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/372.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/373.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_19/Issue_04/tiff/374.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_04/tiff/375.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_04/tiff/376.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/377.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/378.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_19/Issue_04/tiff/379.
tif.

Backus:1961:RAL

[BBG^{+}1 ] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, H. Rutishauser, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language ALGOL 60. Acta Polytechnica Scandinavica. Mathematics and Computing Series, 5:1–39, 1961

(??). ISSN 1456-9418. URL https://web.archive.org/
web/20191029151524id_/http://web.eah-jena.de/~kleine/
history/languages/Algol60-Naur.pdf.

Backus:1958:PPL

[BDE^{+}58] J. W. Backus, P. H. Diselets, D. C. Evans, R. Goodman, H. Huskey, C. Katz, J. McCarthy, A. Orden, A. J. Perlis, R. Rich, S. Rosen, W. Turanski, and J. Wegstein. Proposal for a programming language. ACM Ad Hoc Committee on Languages

report, 1958. URL https://datamuseum.dk/bits/30009315;
 https://datamuseum.dk/wiki/Bits:30009315;   https://
softwarepreservation.computerhistory.org/ALGOL/report/
ACM_ALGOL_Proposal_1958.pdf.

Backus:1954:ISO

[BH54] John W. Backus and Harlan Herrick. IBM 701 Speedcoding and other automatic programming systems. In ????, editor, Proceedings of the Symposium on Automatic Programming for Digital Computers, Navy Mathematical Computing Advisory Panel, 13–14 May 1954, pages 106–113. The Office of Naval Research, Washington, DC, USA, May 1954. URL https://archive.

computerhistory.org/resources/text/Fortran/102653983.
05.01.acc.pdf.

Backus:1964:F

[BH64] J. W. Backus and W. P. Heising. Fortran. IEEE Transactions on Electronic Computers, EC-13(4):382–385, August 1964. CODEN

IEECA8. ISSN 0367-7508. URL https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4038201.

Buchholz:1990:CQDa

[BHW^{+}90] Werner Buchholz, R. Anthony Hyman, Maurice V. Wilkes, Richard E. Smith, George T. Jacobi, Alice R. Burks, Arthur W. Burks, I. Bernard Cohen, I. O. Grattan-Guinness, S. I. Gass, and Mike Woodger. Comments, queries, and debate: Whiggism in the History of Science and the Study of the Life and Work of Charles Babbage; Whirlwind and Microprogramming; More on Wilkes, Whirlwind, and Microprogramming; Pioneer Biographies; The Atanasoff Story — A Response; Notes on Babbage, Aiken, and Bowditch; Joseph Fourier’s Anticipation of Linear Programming; What Does BNF Stand for? Annals of the History of Computing, 12(1):62–72, January/March 1990. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1990/pdf/a1062.pdf;   http://www.computer.org/
annals/an1990/a1062abs.htm.

Bjorner:2008:JWB

[Bjø08] Dines Bjørner. John Warner Backus: 3 Dec 1924–17 March 2007. Formal Aspects of Computing, 20(3):239–240, May 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s00165-008-0077-4;   http://link.springer.
com/article/10.1007/s00165-008-0077-4.

Bruggemann-Klein:1993:UERa

[BK93] A. Brüggemann-Klein. Unambiguity of extended regular expressions in SGML document grammars. In Lengauer [Len93], pages 73–84. ISBN 3-540-57273-2, 0-387-57273-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43E83 1993.

Booch:2006:OHJ

[Boo06] Grady Booch. Oral history of John Backus. Web document.,

September 5, 2006. URL http://archive.computerhistory.
org/resources/text/Oral_History/Backus_John/Backus_John_
1.oral_history.2006.102657970.pdf.

Broy:1985:SAS

[Bro85] Manfred Broy. Structured algebraic specification of Backus’ functional programming language. RAIRO. Technique et Science Informatiques, 4:447–458, 1985. ISSN 0752-4072 (print), 2116-5920 (electronic).

Busa:1962:SLD

[Bus62] Roberto Busa, editor. Symbolic Languages in Data Processing, Proceedings of the Symposium, Rome, March 26–31, 1962. Gordon and Breach Science Publishers, New York, NY, USA, 1962. LCCN QA76 .S95 1962.

Backus:1986:FLM

[BWW86] John W. Backus, J. H. Williams, and E. L. Wimmers. The FL language manual. Technical Report RJ 5339 (54809), IBM Corporation, Armonk, NY, USA, 1986.

Backus:1989:FLM

[BWW^{+}89] J. Backus, J. H. Williams, E. L. Wimmers, P. Lucas, and A. Aiken. FL language manual, parts 1 and 2. Research Report RJ 7100 (67163), IBM Corporation, ????, 1989.

Backus:1990:IPL

[BWW90] J. Backus, J. H. Williams, and E. L. Wimmers. An introduction to the programming language FL. In Turner [Tur90], page ?? ISBN 0-201-17236-4. LCCN QA76.62 .R47 1990.

Cantor:1962:APB

[Can62] David G. Cantor. On the ambiguity problem of Backus systems. Journal of the ACM, 9(4):477–479, October 1962. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chaplin:1973:GRB

[CCH73] R. I. Chaplin, R. E. Crosbie, and J. L. Hay. A graphical representation of the Backus–Naur form. The Computer Journal, 16(1): 28–29, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/16/1/28.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_16/Issue_01/160028.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/28.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_16/Issue_01/tiff/29.
tif.

Chen:2016:GAT

[Che16] Jiahao Chen. “This guy’s arrogance takes your breath away”: Letters between John W Backus and Edsger W Dijkstra, 1979. Web site., May 29, 2016. URL https://medium.com/

@acidflask/this-guys-arrogance-takes-your-breath-away-

5b903624ca5f. See Dijkstra’s response [Dij78] to Backus’ Turing Award Lecture.

Chiarini:1980:BFS

[Chi80] A. Chiarini. BACKUS-FP: un sistema per la programmazione funzionale. Versione 1.0. (Italian) [BACKUS-FP: a system for functional programming. Version 1.0], volume 80. Comitato Nazionale Energia Nucleare, Rome, Italy, 1980.

Cipra:2000:BCE

[Cip00] Barry A. Cipra. The best of the 20th Century: Editors name top 10 algorithms. SIAM News, 33(4):1–2, May 2000. ISSN 0036-1437.

URL https://archive.siam.org/pdf/news/637.pdf.

Campbell-Kelly:2007:JB

[CK07] Martin Campbell-Kelly. John Backus (1924–2007). Nature, 446 (7139):998, April 26, 2007. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Cortada:1987:HDD

[Cor87] James W. Cortada. Historical Dictionary of Data Processing: Biographies. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1987. ISBN 0-313-25651-9. xiii + 321 pp. LCCN QA76.15 .C66 1987.

Culik:1972:BRJb

[Čul72a] Karel Čulík. Book review: J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL 60. Numerische Mathematik, vol. 4 (1962–1963), pp. 420–453; also ibid., vol. 6 (1963), pp. 1–17; and The Computer Journal, vol. 5 no. 4 (1963), pp. 349–367. Journal of Symbolic Logic, 37(3):623– 624, September 1972. CODEN JSYLA6. ISSN 0022-4812 (print),

1943-5886 (electronic). URL https://www.jstor.org/stable/
2272794.

Culik:1972:BRJa

[Čul72b] Karel Čulík. Book review: J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur (editor), A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language ALGOL 60. Numerische Mathematik, vol. 2 (1960), pp. 106–136; also Communications of the Association for Computing Machinery, vol. 3 (1960), pp. 299–314. Journal of Symbolic Logic, 37(3):622–623, September 1972. CODEN JSYLA6. ISSN 0022-4812 (print),

1943-5886 (electronic). URL https://www.jstor.org/stable/
2272793;   ttps://www.jstor.org/stable/2272793.

deBakker:1981:ALP

[dBvV81] J. W. (Jacobus Willem) de Bakker and J. C. van Vliet, editors. Algorithmic languages: Proceedings of the International Symposium on …, Amsterdam, The Netherlands, 26–29 October 1981, a tribute to prof. A. van Wijngaarden on the occasion of his retirement from the Mathematical Centre. North-Holland, Amsterdam, The Netherlands, 1981. ISBN 0-444-86285-4. LCCN QA76.7 .I575

1981. URL https://ir.cwi.nl/pub/34328/34328D.pdf.

Dijkstra:1962:PAR

[Dij62] Edsger W. Dijkstra. A primer of ALGOL 60: report on the algorithmic language ALGOL 60, volume 2 of APIC studies in data processing. Academic Press, New York, USA, 1962. ISBN 0-12-216250-1. ISSN 0067-2483. xi + 114 pp. LCCN QA76.73.A24.

URL https://dn710009.ca.archive.org/0/items/a_primer_
of_algol_60_programming/algol60_o.pdf.

Dijkstra:1978:RTA

[Dij78] Edsger W. Dijkstra. A review of the 1977 Turing Award Lecture by John Backus. Circulated privately., 1978. URL http://www.

cs.utexas.edu/users/EWD/ewd06xx/EWD692.PDF.

Dosch:1983:ASB

[DM83] Walter Dosch and Bernhard Möller. An algebraic semantics for Backus’ functional programming language with infinite objects.

In GI - 13. Jahrestagung, Hamburg, 3–7 Oktober 1983 Proceedings, volume 73, pages 67–85. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983.

DeRemer:1982:ECL

[DP82] Frank DeRemer and Thomas Pennello. Efficient computation of LALR(1) look-ahead sets. ACM Transactions on Programming Languages and Systems, 4(4):615–649, October 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Essalmi:2006:GUV

[EA06] F. Essalmi and L. J. B. Ayed. Graphical UML view from extended Backus–Naur Form grammars. In Kinshuk, Rob Koper, Piet Kommers, Paul Kirschner, Demetrios G. Sampson, and Wim Didderen, editors, Sixth IEEE International Conference on Advanced Learning Technologies (ICALT’06), 5–7 July 2006, Kerkrade, The Netherlands, pages 544–546. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN 0-7695-2632-2. LCCN LB1028.43.I315 2006.

Geppert:1966:BBJ

[Gep66] M. G. Geppert. Buchbesprechung: Backus, J. W., F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. Van Wijngarden und M. Woodger: Bericht über die algorithmische Sprache Algol 60. (Report on the algorithmic language Algol 60. Übers. Kerner J. O. u. Mitarb.). Akademie-Verlag Berlin 1962. vi + 49 S., Preis MDN 10,50. Biometrische Zeitschrift, 8(1-2):121, 1966. CODEN BIZEB3. ISSN 0006-3452.

Gorn:1961:SLM

[Gor61] Saul Gorn. Specification languages for mechanical languages and their processors a baker’s dozen: a set of examples presented to ASA x3.4 subcommittee. Communications of the ACM, 4(12): 532–542, December 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoekstra:2022:CLC

[Hoe22] Conor Hoekstra. Combinatory logic and combinators in array languages. In ARRAY 2022: Proceedings of the 8th ACM SIGPLAN International Workshop on Libraries, Languages and Compilers for Array Programming, San Diego, CA, USA, 13 June 2022, AR-RAY 22, pages 46–57. ACM Press, New York, NY 10036, USA, June 2022. ISBN 1-4503-9269-5. LCCN QA76.76.C65 .A773 2022.

Hudak:1989:CEA

[Hud89] Paul Hudak. Conception, evolution, and application of functional programming languages. ACM Computing Surveys, 21(3): 359–411, September 1989. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL http://www.acm.org/pubs/

toc/Abstracts/0360-0300/72554.html.

Halpern:1987:FAE

[HW87] Joseph Y. Halpern and Edward L. Wimmers. Full abstraction and expressive completeness for FP. In ????, editor, Proceedings of the Second Annual Symposium on Logic in Computer Science (LICS 1987), Ithaca, NY, 22–25 June 1987, pages 246–271. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. ISBN 0-8186-0793-9, 0-8186-4793-0, 0-8186-8793-2. LCCN ???? URL https://lics.siglog.org/

archive/1987/;   https://lics.siglog.org/archive/1987/
HalpernWimmers-FullAbstractionandE.html.

Halpern:1988:SSC

[HW88] Joseph Y. Halpern and Edward L. Wimmers. Sacrificing simplicity for convenience: Where do you draw the line? In J. Ferrante and P. Mager, editors, Proceedings of the 15th ACM SIGPLAN– SIGACT Symposium on Principles of Programming Languages (POPL ’88), San Diego, California, USA, January 10–13, 1988, pages 169–179. ACM Press, New York, NY 10036, USA, 1988. ISBN 0-89791-252-7. LCCN QA76.7 .A15 1988. URL https://

dl.acm.org/doi/10.1145/73560.73575.

Halpern:1995:FAE

[HW95] Joseph Y. Halpern and Edward L. Wimmers. Full abstraction and expressive completeness for FP. Information and Computation, 118(2):246–271, May 1, 1995. CODEN INFCEC. ISSN 0890-5401 (print), 1090-2651 (electronic).

Halpern:1986:GRS

[HWW86] Joseph Y. Halpern, John H. Williams, and Edward L. Wimmers. Good rewrite strategies for FP. In ????, editor, Proceedings of the First Annual IEEE Symposium on Logic in Computer Science (LICS 1986), Cambridge, MA, USA, 16–18 June 1986, pages 149–162. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring,

MD 20910, USA, 1986. URL https://lics.siglog.org/
archive/1986/;   https://lics.siglog.org/archive/1986/
HalpernWilliamsWimm-GoodRewriteStrategi.html.

Halpern:1990:CRR

[HWW90] Joseph Y. Halpern, John H. Williams, and Edward L. Wimmers. Completeness of rewrite rules and rewrite strategies for FP. Journal of the ACM, 37(1):86–143, January 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/78939.html.

Halpern:1985:DSR

[HWWW85] Joseph Y. Halpern, John H. Williams, Edward L. Wimmers, and Timothy C. Winkler. Denotational semantics and rewrite rules for FP. In Mary S. Van Deusen, Zvi Galil, and Brian K. Reid, editors, Proceedings of the 12th ACM SIGACT-SIGPLAN symposium on Principles of programming languages — POPL ’85, New Orleans, Louisiana, USA, January 14–16, 1985, POPL ’85, pages 108–120. ACM Press, New York, NY 10036, USA, 1985. ISBN 0-89791-147-4.

IBM:1953:ISS

[IBM53] IBM. IBM Speedcoding System for the [IBM] Type 701. IBM Corporation, New York, NY, USA, September 1953. Likely written by John Backus and his team.

IBM:1954:SIM

[IBM54] IBM. Specifications for the IBM Mathematical FORmula TRANSlating system. Preliminary report, IBM Corp., Programming Research Group, Applied Sciences Division, New York, NY, USA,

November 10, 1954. URL https://archive.computerhistory.
org/resources/text/Fortran/102679231.05.01.acc.pdf;
https://softwarepreservation.computerhistory.org/FORTRAN/
BackusEtAl-Preliminary%20Report-1954.pdf.

IBM:20xx:JB

[IBMxx] IBM. John Backus. Web site., 20xx. URL http://www-03.

ibm.com/ibm/history/exhibits/builders/builders_backus.

html. Includes an interview.

ISO:1991:ISI

[II91] International Organization for Standardization and International Electrotechnical Commission. International standard: information, technology, programming languages, Fortran. International Organization for Standardization, Geneva, Switzerland, second edition, 1991. ISBN ???? xvii + 369 pp. LCCN ???? URL http://

www.iso.ch/cate/d26933.html;   http://www.iso.ch/cate/
d26934.html; http://www.iso.ch/cate/d29926.html.

Ingerman:1962:TTL

[Ing62] P. Z. Ingerman. A translation technique for languages whose syntax is expressible in Backus Normal Form. In Busa [Bus62], pages 23–64. LCCN QA76 .S95 1962.

Ingerman:1967:PBF

[Ing67] Peter Zilahy Ingerman. “Pānini–Backus Form” suggested. Communications of the ACM, 10(3):137, March 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/pdf/10.1145/363162.363165.

IBM:1957:TYF

[Int57] International Business Machines Corporation. Twenty-five years of FORTRAN exhibit, 1957.

IBM:1959:POM

[Int59] International Business Machines Corporation. Preliminary Operator’s Manual: Automatic Coding System for IBM 704 EDPM (Fortran 4-1-4-1). IBM Corporation, New York, NY, USA, April

8, 1959. 37 pp. URL https://archive.computerhistory.
org/resources/text/Fortran/102679276.05.01.acc.pdf;
https://softwarepreservation.computerhistory.org/FORTRAN/
manual/Prelim_Oper_Man-1957_04_07.pdf. Includes copy of

[BBB^{+}57], plus two letters by J. Backus.

ISO:1994:IIIe

[ISO94] ISO. ISO/IEC 1539-2:1994: Information technology — Programming languages — FORTRAN — Part 2: Varying length character strings. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 66 pp. LCCN ????

CHF 152. URL http://www.ansi.org/;   http://www.iso.

ch/cate/d6129.html. Available in English only.

ISO:1997:IIIa

[ISO97a] ISO. ISO/IEC 1539-1:1997: Information technology — Programming languages — Fortran — Part 1: Base language. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? ???? pp. LCCN ???? US$195.00.

URL http://webstore.ansi.org/ansidocstore/product.asp?
sku=ISO%2FIEC+1539%2D1%3A1997.

ISO:1997:IIIg

[ISO97b] ISO. ISO/IEC 1539-1:1997: Information technology — Programming languages — Fortran — Part 1: Base language. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 346 pp. LCCN ???? CHF 288. URL http://

www.ansi.org/;   http://www.fortran.com/fortran/Books/
f95std.html;   http://www.iso.ch/cate/d26933.html. ISO/IEC

JTC 1/SC 22/WG 5. This is the Fortran 95 Standard. Available in English only.

ISO:1998:IITc

[ISO98a] ISO. ISO/IEC TR 15580:1998: Information technology — Programming languages — Fortran — Floating-point exception handling. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 27 pp. LCCN ???? CHF 104;

US$72.00. URL http://webstore.ansi.org/ansidocstore/
product.asp?sku=ISO%2FIEC+TR+15580%3A1998;   http://www.

iso.ch/cate/d28230.html. Available in English only.

ISO:1998:IITd

[ISO98b] ISO. ISO/IEC TR 15581:1998: Information technology — Programming languages — Fortran — Enhanced data type facilities. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 13 pp. LCCN ???? CHF 68; US$46.00.

URL http://webstore.ansi.org/ansidocstore/product.asp?
sku=ISO%2FIEC+TR+15581%3A1998;   http://www.iso.ch/cate/

d28231.html. Available in English only.

ISO:1999:IIIe

[ISO99] ISO. ISO/IEC 1539-3:1999: Information technology — Programming languages — Fortran — Part 3: Conditional compilation. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 21 pp. LCCN ???? CHF 92; US$62.00.

URL http://webstore.ansi.org/ansidocstore/product.asp?
sku=ISO%2FIEC+1539%2D3%3A1999;   http://www.iso.ch/cate/

d29926.html. Available in English only.

ISO:2000:IIIe

[ISO00] ISO. ISO/IEC 1539-2:2000: Information technology — Programming languages — Fortran — Part 2: Varying length character strings. International Organization for Standardization, Geneva, Switzerland, 2000. ISBN ???? ???? pp. LCCN ???? US$58.00.

URL http://webstore.ansi.org/ansidocstore/product.asp?
sku=ISO%2FIEC+1539%2D2%3A2000.

ISO:2004:DIS

[ISO04a] ISO. Draft International Standard ISO/IEC 1539-1:2004(E): Information technology — Programming languages — Fortran Part 1: Base Language. International Organization for Standardization, Geneva, Switzerland, May 2004. xiv + 569 pp. URL ftp:/

/ftp.nag.co.uk/sc22wg5/N1601-N1650/N1601.pdf.gz.

ISO:2004:IIIa

[ISO04b] ISO. ISO/IEC 1539-1:2004 Information technology — Programming languages — Fortran – Part 1: Base language. International Organization for Standardization, Geneva, Switzerland, 2004. xiv

+ 569 pp. URL http://www.dkuug.dk/jtc1/sc22/open/n3661.
pdf.

ISO:2010:IIIb

[ISO10] ISO. ISO/IEC 1539-1:2010 Information technology — Programming languages — Fortran — Part 1: Base language. International Organization for Standardization, Geneva, Switzerland,

June 2010. xviii + 603 pp. URL ftp://ftp.nag.co.uk/sc22wg5/
N1801-N1850/N1830.pdf.

Johansen:1966:CRD

[Joh66]  Peter Johansen.  Construction of recognition devices for

regular languages from their Backus Normal Form definition. Nordisk tidskrift for informationsbehandling, 6(4):294– 309, July 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0006-3835&volume=6&
issue=4&spage=294.

Knuth:1962:BL

[KG62] Donald E. Knuth and Saul Gorn. Letters to the Editor: Backus’ language. Communications of the ACM, 5(4):185, April 1962. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gor61].

Kusmierz:1974:AMP

[KJ74] Grazyna Kusmierz and Ryszard Janda. Application method of program-generator of any language written in Backus notation as an acceptor of this language. Algorytmy, 11(20):85–135, 1974.

Knuth:1964:LEB

[Knu64] Donald E. Knuth. Letter to the editor: Backus Normal Form vs. Backus Naur Form. Communications of the ACM, 7(12):735– 736, December 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Knuth:2003:SPC

[Knu03] Donald E. Knuth. Selected Papers on Computer Languages, volume 139 of CSLI Lecture Notes. CSLI Publications, Stanford, CA, USA, 2003. ISBN 1-57586-381-2 (hardcover), 1-57586-382-0 (paperback). xviii + 594 pp. LCCN QA76.7 .K63 2002.

Konstantinov:1973:CTG

[Kon73] V. I. Konstantinov. Konstruktion von TR-Grammatiken nach Grammatiken in der Backus–Naur-Form. (German) [Construction of TR-grammars based on grammars in the Backus–Naur form]. Vycislit. Sistemy, Novosibirsk, 57:91–97, 162, 1973. ISSN 0568-661x.

Laplante:1996:GPC

[Lap96] Phillip Laplante, editor. Great Papers in Computer Science. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-314-06365-X (paperback), 0-7803-1112-4 (hardcover). iv + 717 pp. LCCN QA76 .G686 1996.

URL http://bit.csc.lsu.edu/~chen/GreatPapers.html. Con-

tains a reprint of the first published paper on Fortran [BBB^{+}57].

LEcuyer:1987:FFR

[L’E87] Pierre L’Ecuyer. Formal formatting rules for Pascal programs. The Journal of Systems and Software, 7(4):311–322, December 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lengauer:1993:AEF

[Len93] T. Lengauer, editor. Algorithms — ESA’93. First Annual European Symposium Proceedings: September 1993, Bad Honnef, Germany, volume 726 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-57273-2, 0-387-57273-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43E83 1993.

Lewin:1994:FDR

[Lew94] David I. Lewin. Fortran’s developer receives top engineering award. Computers in Physics, 8(1):6–7, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208

(electronic). URL https://aip.scitation.org/doi/10.1063/
1.4823261.

Lohr:2007:JWB

[Loh07] Steve Lohr. John W. Backus, 82, Fortran developer, dies. New York Times obituary., March 19, 2007. URL http:

//www.nytimes.com/2007/03/19/obituaries/20cnd-backus.
html?ex=3D1332043200&en=3Dadde3ee5a1875330&ei=3D5124&
partner=3Dpermalink&exprod=3Dpermalink.

Lucas:1981:FSP

[Luc81] P. Lucas. Formal semantics of programming languages: VDL. IBM Journal of Research and Development, 25(5):549–561, September 1981. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Laning:1954:PTM

[LZ54] J. Halcombe Laning, Jr. and Neal Zierler. A program for translation of mathematical equations for Whirlwind I. Engineering Memorandum E-364, Instrumentation Laboratory, MIT, Cambridge, MA, USA, January 1954.

McDonald:1989:DSC

[MA89] C. S. McDonald and L. Allison. Denotational semantics of a command interpreter and their implementation in Standard ML. The Computer Journal, 32(5):422–431, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/5/422.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_32/Issue_05/tiff/422.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_32/
Issue_05/tiff/423.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/424.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
05/tiff/425.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_05/tiff/426.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/427.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/428.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/429.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
05/tiff/430.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_05/tiff/431.tif.

McJones:1975:CRP

[McJ75] Paul McJones. A Church–Rosser property of closed applicative languages. Research Report RJ 1589, IBM Research Laboratory, San Jose, CA, USA, May 1975. URL http://www.mcjones.org/

paul/rj1589.pdf.

McJones:2017:RJB

[McJ17] Paul McJones. Remembering John Backus. Web site., April 1,

2017. URL https://www.mcjones.org/dustydecks/archives/
2007/04/01/60/.

Meyer:1990:ITP

[Mey90] Bertrand Meyer. Introduction to the Theory of Programming Languages. Prentice Hall international series in computer sciences. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-498502-8, 0-13-498510-9. xvi + 447 pp. LCCN QA76.7 .M49

1988. URL https://bertrandmeyer.com/ITPL.

Metropolis:1980:HCT

[MHR80] Nicholas Metropolis, Jack Howlett, and Gian-Carlo Rota, editors. A History of Computing in the Twentieth Century: a Collection of Essays. Academic Press, New York, USA, 1980. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Malita:2011:BLF

[MŞ11] Mihaela Maliţa and Gheorghe M. Ştefan. Backus language for functional nano-devices. In CAS 2011 Proceedings (2011 International Semiconductor Conference), volume 2, pages 331–334. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011.

Naur:1964:RRA

[Nau64] Peter Naur, editor. Revised report on the algorithmic language ALGOL 60. Regnecentralen, Copenhagen, Denmark, 1964. 43 pp. LCCN QA76.5 .N35 1964. Written by John W. Backus and others.

Naur:1960:RALa

[NBB^{+}60] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language ALGOL 60. Communications of the ACM, 3(5):299–314, May 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://web.

archive.org/web/20171002221007id_/http://www.eah-jena.
de/~kleine/history/languages/Algol60-ACM.pdf. See letter

[AGG60]. Dedicated to the Memory of William Turanski [American group member who was killed by an automobile just prior to the January 1960 Conference].

Naur:1962:BAS

[NBB^{+}62] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, A. J. Perlis, S. Ruti shauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Bericht über die algorithmische Sprache ALGOL 60. (German) [Report on the ALGOL 60 language]. Elektronisches Rechnen und Regeln. Sonderband 1. Akademie-Verlag, Berlin, Germany, 1962. xi + 49 pp. LCCN QA76.5 .N3515. Zur Erinnerung an William Turanski (In memory of William Turanski).

Naur:1963:RRA

[NBB^{+}63] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, M. Woodger, and P. Nauer. Revised report on the algorithmic language ALGOL 60. Communications of the ACM, 6(1):1–17, January 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Dedicated to the memory of William Turanski.

Naur:1965:RRA

[NBB^{+}65] P. Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, Heinz Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL-60. (Russian), volume Moskau, Russia. Verlag “Mir”, 1965. 80 pp.

Naur:1997:RRA

[NBB^{+}97] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Kafz, J. Mc-Carthy, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language Algol 60. In Algol-like Languages, pages 19–49. Birkhäuser Boston Inc., Cambridge, MA, USA, 1997.

Nechesov:2022:SQPb

[Nec22a] A. V. Nechesov. Some questions on polynomially computable representations for generating grammars and Backus–Naur forms. Siberian Advances in Mathematics, 32(4):299–309, 2022. ISSN 1055-1344 (print), 1934-8126 (electronic). Translation of Matematicheskie Trudy 25 (2022), no. 1, 134–151.

Nechesov:2022:SQPa

[Nec22b] Andreĭ Vital’evich Nechësov. Some questions on polynomially computable representations for generating grammars and Backus– Naur forms. Matematicheskie Trudy, 25(1):134–151, 2022. ISSN 1560-750X.

Nuding:1970:COR

[NKW70] E. Nuding and I. Kahlert-Warmbold. A computer-oriented representation of matrices. Computing, 6(1–2):1–8, March 1970. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nori:1986:GBF

[Nor86] Kesav V. Nori. A generalization of Backus’ FP. In Srinivas and Sangal [SS86], pages 124–143. ISBN 0-387-17179-7, 3-540-17179-7, 3-540-47239-8. LCCN QA76.751 .F68 1986.

Pagan:1986:FTB

[Pag86] Frank G. Pagan. On the feasibility of teaching Backus-type functional programming (FP) as a first language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 18(3):31–35, September 1, 1986. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Petrick:1965:MBN

[Pet65] S. R. Petrick. More on Backus Normal Form. Communications of the ACM, 8(3):200–201, March 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Letter to the editor.

Post:1943:FRG

[Pos43] Emil L. Post. Formal reductions of the general combinatorial decision problem. American Journal of Mathematics, 65:197–215, 1943. CODEN AJMAAN. ISSN 0002-9327,1080-6377.

Popov:2024:ABN

[PP24] George Popov and Antoaneta Popova. Application of Backus– Naur Forms for investigation [of] financial customs violations. In ????, editor, 2024 12th International Scientific Conference on Computer Science (COMSCI), pages 1–4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2024.

Radensky:1987:LEN

[Rad87] A. Radensky. Lazy evaluation and nondeterminism make Backus’ FP-systems more practical. ACM SIGPLAN Notices, 22(4):33–40, April 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Randell:1991:PBB

[Ran91] Brian Randell. Programming: from Babbage to Backus. Technical report 327, University of Newcastle upon Tyne, Computing Laboratory, Newcastle upon Tyne, UK, June 1991. 17 pp.

URL https://eprints.ncl.ac.uk/file_store/production/
160290/DA7EC111-C6F7-4AB1-8820-98112C4BA279.pdf.

Radensky:1989:SRE

[RD89] Atanas Radensky and Milena Djambazova. Schemes for recursion elimination in Backus’ FP-systems. Godishnik na Sofiĭskiya Universitet “Sv. Kliment Okhridski”. Fakultet po Matematika i Informatika, 83(1):39–49, 1989. ISSN 0205-0811.

Revesz:1992:LOE

[Rev92] György E. Revesz. A list-oriented extension of the lambdacalculus satisfying the Church–Rosser theorem. Theoretical Computer Science, 93(1):75–89, February 03, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
030439759290212X.

Robinet:1980:LBS

[Rob80] B. Robinet. Les langages de Backus sont des systèmes de manipulation d’arbres. (French) [Backus’ languages are tree-manipulation systems]. In M. Dauchet and E. Lilin, editors, Les arbres en algèbre et en programmation, 5ième Colloque, Lille, 23–25 fèvrier 1980, pages 83–94. Université de Lille, Lille, France, 1980.

Robison:1987:IFPb

[Rob87a] A. D. Robison. The Illinois functional programming interpreter. In Richard L. Wexelblat, editor, Papers of the Symposium on Interpreters and interpretive techniques — SIGPLAN ’87, St. Paul, Minnesota, USA, June 24–26, 1987, SIGPLAN 87, pages 64–73. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-235-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7 A84 v. 22, no. 7.

Robison:1987:IFPa

[Rob87b] Arch D. Robison. Illinois functional programming: a tutorial. Byte Magazine, 12(2):115–125, February 1987. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (elec-

tronic). URL https://archive.org/details/BYTE-MAGAZINE-
COMPLETE/198702_Byte_Magazine_Vol_12-02_Image_Processing/
page/n139.

Rohl:1968:NBN

[Roh68] J. S. Rohl. A note on Backus Naur Form. The Computer Journal, 10(4):336–337, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/10/4/336.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_04/100336.sgm.abs.html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_04/tiff/336.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_04/tiff/337.tif.

Rosen:1966:PSL

[Ros66] Saul Rosen, editor. Programming Systems and Languages. McGraw-Hill computer science series. McGraw-Hill, New York, NY, USA, 1966. xv + 734 pp. LCCN QA76.5 .R53. Reissued in 1967. Contains a reprint of the first published paper on Fortran [BBB^{+}57] on pp. 29–47.

Sammet:1981:HIT

[Sam81] Jean E. Sammet. History of IBM’s technical contributions to high level programming languages. IBM Journal of Research and Development, 25(5):520–534, September 1981. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Saphire:1967:IJB

[Sap67] Larry Saphire. Interview of John Backus, San Francisco, California. IBM Oral History of Computer Technology TC-46, IBM Corporation, New York, NY, USA, December 15, 1967. 55

pp. URL https://softwarepreservation.computerhistory.
org/FORTRAN/paper/Saphire_Interview_of_John_Backus-TC-
46-1967_12_15.pdf.

Sheridan:1959:ATC

[She59] Peter B. Sheridan. The arithmetic translator compiler of the IBM FORTRAN Automatic Coding System. Communications of the ACM, 2(2):9–21, February 1959. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Steele:1993:FCE

[SHS^{+}93] Guy L. Steele, Jr., Danny Hillis, Richard Stallman, Gerald J. Sussman, Marvin Minsky, John McCarthy, John Backus, Fernando Corbató, and Ronald E. Anderson. Forum: Code of ethics reconsidered. Communications of the ACM, 36(7):17–20, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shasha:1998:TML

[SL98] Dennis Shasha and Cathy Lazere, editors. Out of Their Minds: The Lives and Discoveries of 15 Great Computer Scientists. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98269-8. xi + 293 pp. LCCN QA76.2.A2 S53 1998. US$16.00.

Slater:1987:JBM

[Sla87a] Robert Slater. John Backus: The man who invented FOR-TRAN. In Portraits in Silicon [Sla87b], chapter 21, pages 230– 239. ISBN 0-262-19262-4 (hardcover), 0-262-69131-0 (paperback), 0-262-29142-8 (e-book). LCCN TK7885.2 .S571 1987.

Slater:1987:PS

[Sla87b] Robert Slater. Portraits in Silicon. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-19262-4 (hardcover), 0-262-69131-0 (paperback), 0-262-29142-8 (e-book). xiv + 374 pp. LCCN TK7885.2 .S571 1987.

Spector:1983:LPM

[Spe83] David Spector. Lexing and parsing Modula-2. ACM SIGPLAN Notices, 18(10):25–32, October 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Srinivas:1986:GBF

[SS86] Y. V. Srinivas and Rajeev Sangal, editors. Foundations of software technology and theoretical computer science, Proceedings of the 6th Conference, New Delhi, India December 18–20, 1986, volume 241 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-17179-7, 3-540-17179-7, 3-540-47239-8. LCCN QA76.751 .F68 1986.

Stegmann:1979:PYP

[Ste79] Claire Stegmann. Pathfinder: For 25 years a programming pioneer John Backus still crackles with ideas. Think Magazine, 45(4):18–

24, July/August 1979. URL https://lc.fie.umich.mx/~legg/
clases/paradigmasdeprogramacion/Backus-Think.pdf.

Torii:1987:LPT

[TSF^{+}87] Koji Torii, Yuji Sugiyama, Mamoru Fujii, Tadao Kasami, and Yoshitomi Morisawa. Logical programming for the telegram analysis problem. Computer Languages, 12(1):9–20, ???? 1987. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Turner:1990:RTF

[Tur90] David A. Turner, editor. Research topics in functional programming: Papers from a meeting held at the University of Texas at Austin, August 1987, University of Texas at Austin Year of Programming series. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-17236-4. LCCN QA76.62 .R47 1990.

Weiss:1988:BOP

[Wei88] Eric A. Weiss. Biographies: Oh, pioneers! Annals of the History of Computing, 10(4):348–361, October/December 1988. CODEN

AHCOE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1988/pdf/a4348.pdf;   http://www.computer.
org/annals/an1988/a4348abs.htm.

Wexelblat:1981:HPL

[Wex81] Richard L. Wexelblat, editor. History of programming languages. Proceedings of the ACM SIGPLAN conference (Los Angeles, Calif., June 1–3, 1978). Academic Press, New York, USA, June 1981. ISBN 0-12-745040-8. LCCN QA76.7 .H56 1978. URL

https://dl.acm.org/doi/book/10.1145/800025.

Williams:1980:DAF

[Wil80] John H. Williams. On the development of the algebra of functional programs. Report RJ2983, IBM Research Laboratory, San Jose, CA, USA, 1980.

Williams:1981:NFS

[Wil81] John H. Williams. Notes on the FP style of functional programming. Lecture notes for the course Functional Programming and its Applications, University of Newcastle upon Tyne., July 1981.

Williams:1982:NFS

[Wil82a] John H. Williams. Notes on the FP style of functional programming. In J. Darlington, P. Henderson, and D. A. Turner, editors, Functional Programming and its Applications, page ?? Cambridge University Press, Cambridge, UK, 1982.

Williams:1982:DAF

[Wil82b] John H. Williams. On the development of the algebra of functional programs. ACM Transactions on Programming Languages and Systems, 4(4):733–757, October 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Y:1977:JBF

[Y.77] E. K. Y. John Backus of Fortran fame: how he’d unclog the ‘von Neumann’ bottleneck. Datamation, 23(1):142–143, 145, January 1977. CODEN DTMNAT. ISSN 0011-6963. URL https://www.

bitsavers.org/magazines/Datamation/197701.pdf.